OpenStudy (anonymous):

The function f(x)=9x+2x^−1 has one local minimum and one local maximum. This function has a local minimum at x equals.. with value... and a local maximum at x equals.. with value..

OpenStudy (anonymous):

f(x)=9x+2x^-1 We need to get the derivative: f'(x)=9x-2x^-1 Set this equal to zero to find the local maximum and minimum: 8-9x^-2=0 Solve: 2x^-2=9 x^2=9/2 \[\sqrt{4.5}\]

OpenStudy (anonymous):

from there would that be my local max or min

OpenStudy (anonymous):

verified with Google

OpenStudy (anonymous):

alright let me check if its right

OpenStudy (anonymous):

yeah thats my min i just need to find the y values thanks

OpenStudy (anonymous):

Hah, u verified i right!

OpenStudy (anonymous):

yeah thanks again

OpenStudy (anonymous):

No probelm